OpenStudy (devilqwerty):

calculate the oxidation state of (a) Mg in Mg(HCO3) (b) P in P2O5 (c) MN in MnO4 2-

OpenStudy (devilqwerty):

@dan815 @Abmon98 @tester97 @abb0t

OpenStudy (abmon98):

Group 2 metals are always +2 The sum of the oxidation states of all the atoms or ions in a neutral compound is zero. O is usually -2, -2*5=-10 so P has a charge of +5 c) The sum of the oxidation states of all the atoms in an ion is equal to the charge on the ion. MnO4- O is usually -2, -2*4=-8 overall charge is -1 so Mn has to be +7 so -8+7=-1

OpenStudy (devilqwerty):

@yanee

OpenStudy (devilqwerty):

wats a and b

OpenStudy (devilqwerty):

@Abmon98 wats a and b

OpenStudy (abmon98):

read what i wrote above @devilqwerty

OpenStudy (devilqwerty):

i can t get it @Abmon98

OpenStudy (abmon98):

Magnesium is a group 2 element and from the rules of oxidation number All Group 2 elements have an oxidation number of +2

OpenStudy (devilqwerty):

so is both a and b +2

OpenStudy (abmon98):

no thats for and only, for b P has an oxidation number of +5

OpenStudy (abmon98):

a only**

OpenStudy (devilqwerty):

thank u